OpenStudy (anonymous):

Solve for x -5|x+1|=10

OpenStudy (anonymous):

start with \[5|x+10|=x-10\]

OpenStudy (anonymous):

then break in to two cases if \(x\geq -1\) then \(|x+1|=x+1\) in that case, solve \[5(x+1)=x-10\]

OpenStudy (anonymous):

oh i made a typo above, sorry should be \[5|x+1|=x-10\]

OpenStudy (anonymous):

Okay, but what would x be?